# PCF Learned Sort: a Learning Augmented Sort Algorithm with $\mathcal{O}(n\log\log n)$ Expected Complexity

### v1

@@ Abstract
Sorting is one of the most fundamental algorithms in computer science. Recently, Learned Sorts, which use machine learning to improve sorting speed, have attracted attention. While existing studies show that Learned Sort is experimentally faster than classical sorting algorithms, they do not provide theoretical guarantees about its computational complexity. We propose PCF Learned Sort, a theoretically guaranteed Learned Sort algorithm. We prove that the expected complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$ under mild assumptions on the data distribution. We also confirm experimentally that PCF Learned Sort has a computational complexity of $\mathcal{O}(n\log\log n)$ on both synthetic and real datasets. This is the first study to theoretically support the experimental success of Learned Sort, and provides evidence for why Learned Sort is fast.
sorting, learning-augmented algorithms, algorithms with predictions
[…]
do
@@ 1 Introduction
The sorting algorithm is one of the most important computer algorithms and has been studied for a long time. Recently, a novel sorting method called Learned Sort has been proposed (Kraska et al., 2019). Learned Sort uses a model that predicts the cumulative density function (CDF) to predict the position of each element in the sorted array, and then uses this prediction to quickly sort the array. Experimental results show that Learned Sort is faster than classical sorting algorithms, including highly tuned counting-based sorting algorithms, comparison sorting algorithms, and hybrid sorting algorithms.
On the other hand, there are few theoretical guarantees about the computational complexity of Learned Sort. The first proposed Learned Sort algorithm (Kraska et al., 2019) has a best-case complexity of $\mathcal{O}(n)$, but its expected or worst-case complexity isn’t discussed. The more efficient Learned Sort algorithms proposed later (Kristo et al., 2020, 2021) also have $\mathcal{O}(n)$ best-case complexity, but $\mathcal{O}(n^{2})$ worst-case complexity (or $\mathcal{O}(n\log n)$ with some modifications). The goal of this paper is to provide a Learned Sort that is theoretically guaranteed to be computationally efficient.
We propose PCF Learned Sort, which can sort with expected complexity $\mathcal{O}(n\log\log n)$ under mild assumptions on the data distribution. Furthermore, we prove that PCF Learned Sort can sort with worst-case computational complexity $\mathcal{O}(n\log n)$. Then, we experimentally confirmed that our Learned Sort can sort with the average complexity of $\mathcal{O}(n\log\log n)$ on both synthetic and real datasets.
In Section 2, we first present related work. Then, in Section 3, we introduce our PCF Learned Sort and give the theorems and brief explanations of their proofs that guarantee its computational complexity. Then, in Section 4, we verify our theorems experimentally. We then discuss the limitations and possible future work in Section 5. Finally, in Section 6, we summarize our results and conclusions.
@@ 2 Related Work
Our research is in the context of algorithms with predictions, in particular, the sorting algorithms with predictions (Section 2.1). There are two types of sorting algorithms, comparison sorts (Section 2.2) and distribution sorts (Section 2.3), and our proposed method is a distribution sort. However, the idea, implementation, and proof of computational complexity of our method are similar to those of sample sort, which is a type of comparison sort. Furthermore, our proposed algorithm and proof of computational complexity are based on those of Learned Index (Section 2.4).
2.1 Algorithms with predictions
Our research is in the area of algorithms with predictions (Mitzenmacher & Vassilvitskii, 2022). These studies focus on improving the performance (e.g., speed and memory usage) of algorithms for problems such as ski rental (Purohit et al., 2018; Gollapudi & Panigrahi, 2019; Shin et al., 2023), caching (Narayanan et al., 2018; Rohatgi, 2020; Lykouris & Vassilvitskii, 2021; Im et al., 2022), scheduling (Gollapudi & Panigrahi, 2019; Lattanzi et al., 2020; Lassota et al., 2023), and matching (Antoniadis et al., 2023; Dinitz et al., 2021; Sakaue & Oki, 2022) by using predictions from oracles such as machine learning models. The main idea of these novel algorithms is to exploit the characteristics of the data by actively exploiting information not explicitly used by classical algorithms.
In particular, our work is based on research about sorting with predictions. (Kraska et al., 2019) proposed a method for fast sorting that uses a model $\tilde{F}(q)$ trained to approximate the cumulative distribution function (CDF) $F(q)$. This first Learned Sort roughly sorts the array by placing each element $x$ of the array on the $nF(x)$-th entry of the output array. It then sorts the array completely by refining the roughly sorted array using insertion sort. This algorithm has a best case complexity of $\mathcal{O}(n)$, but a worst case complexity of $\mathcal{O}(n^{2})$ because of the exception handling for collisions and the last refining.
Subsequently, implementations have been proposed that improve cache efficiency (Kristo et al., 2020) and are robust against key duplication (Kristo et al., 2021). These also use insertion sorting in the last refinement, resulting in a worst-case complexity of $\mathcal{O}(n^{2})$. The worst-case complexity can be reduced to $\mathcal{O}(n\log n)$ by using a sorting algorithm with small worst-case complexity, such as Introsort (Musser, 1997) or TimSort (McIlroy, 1993). However, this is the same complexity as many effective classical sorting algorithms, and thus does not explain the experimental success of Learned Sort. Some studies have focused on efficiently exploiting the predictions of the oracle, considering the oracle as a black box and not focusing on its inner details (Bai & Coester, 2023). This analysis does not take into account the computational complexity of training and inferring the oracle, so it cannot be directly applied to guarantee the overall computational complexity of sorting.
@@ 2.2 Comparison sorts
Sorting algorithms that use comparisons between keys and require no other information about the keys are called comparison sorts. It can be proved that the worst complexity of a comparison sort is at least $\mathcal{O}(n\log n)$. Commonly used comparison sorting algorithms include quick sort, heap sort, merge sort, and insertion sort. The GNU Standard Template Library in C++ uses Introsort (Musser, 1997), an algorithm that combines quick sort, heap sort, and insertion sort. Java (Java, 2023) and Python up to version 3.10 (Peters, 2002) use TimSort (McIlroy, 1993), an improved version of merge sort. Python 3.11 and later use Powersort (Munro & Wild, 2018), a merge sort that determines a near-optimal merge order.
Sample sort (Frazer & McKellar, 1970) is an extension of quick sort that uses multiple pivots, whereas quick sort uses only one pivot. Sample sort samples a small number of keys from the array, determines multiple pivots, and uses them to partition the array into multiple buckets. The partitioning is repeated recursively until the array is sufficiently small. The in-place parallel superscalar sample sort (Axtmann et al., 2022) is one of the most efficient sample sort implementation. They theoretically guarantee its computational and cache efficiency by proving a theorem about the probability that the pivots partition the array (nearly) equally.
2.3 Distribution sorts
Distribution sorts use information other than key comparison. Radix sort and counting sort are the most common types of distribution sorts. Radix sort uses counting sort as a subroutine for each digit. When the number of digits in the array element is $w$, the computational complexity of radix sort is $\mathcal{O}(wn)$. Thus, radix sort is particularly effective when the number of digits is small. There are several variants of radix sort, such as SkaSort (Skarupke, 2016), which is optimized using a technique called “American flag sort,” and RegionSort (Obeya et al., 2019), which enables efficient parallelization by modeling and resolving dependencies among the element swaps.
@@ 2.4 Learned Index
Kraska et al. showed that index data structures such as B-trees and Bloom filters can be made faster or more memory efficient by combining them with machine learning models, and named such novel data structures Learned Index (Kraska et al., 2018). Since then, various learning augmented B-trees (Wang et al., 2020; Kipf et al., 2020) and learning augmented Bloom filters (Mitzenmacher, 2018; Vaidya et al., 2021; Sato & Matsui, 2023) have been proposed. There are several researches on learning augmented B-trees whose performance is theoretically guaranteed. PGM-index (Ferragina & Vinciguerra, 2020) is a learning augmented B-tree that is guaranteed to have the same worst-case query complexity as the classical B-tree, i.e., $\mathcal{O}(\log n)$. (Zeighami & Shahabi, 2023) proposed a learning augmented B-tree with an expected query complexity of $\mathcal{O}(\log\log n)$ under mild assumptions on the distribution.
@@ 3 Methods
Figure 1: PCF Learned Sort: First, the input array is partitioned into $\gamma+1$ buckets using a CDF model-based method. Buckets larger than $\delta$ or smaller than $\tau$ are sorted with a “standard” sorting algorithm of complexity $\mathcal{O}(n\log n)$ (e.g. IntroSort). Otherwise, the recursive model-based bucketing is repeated. Finally, the sorted arrays are concatenated. The CDF model used for bucketing is a Piecewise Constant Function (PCF). The function is constant within each interval and the interval widths are constant.
@@
[…]
This section describes our PCF Learned Sort. First, in Section 3.1, we give an overview of the algorithm and the key theorems used to guarantee the expected and worst-case computational complexity. Next, in Section 3.2, we introduce the specific algorithm for PCF Learned Sort and give a guarantee on its computational complexity.
3.1 Method Overview
Our algorithm basically repeats recursive model-based bucketing until the array is small enough or the bucketing “fails.” In model-based bucketing, a CDF model is trained and then used to partition an array of length $n$ into approximately $\gamma$ buckets, satisfying the following two conditions: the elements in $i$-th bucket are smaller than any elements in $(i+1)$-th bucket, and the buckets have approximately the same size with high probability. If we set $\gamma=n^{c}$ (where $10<c<$ 1 is a constant) and the partitioning into buckets is done well, the size of each bucket will be approximately $n^{1-c}$, so the size of the bucket in the $k$-th recurrence depth will be approximately $n^{(1-c)^{k}}$. Here,
$k\geq-\frac{\log\log n}{\log(1-c)}\Rightarrow n^{(1-c)^{k}}$ $\leq n^{(1-c)^{-\frac{\log\log n}{\log(1-c)}}}$ (1)
$=n^{\frac{1}{\log n}}$ (2)
$=e,$ (3)
where $\log$ is the natural logarithm, so the base is $e$. That is, with a recursion depth of $k=\Omega(\log\log n)$, the size of bucket is small enough, i.e., the max recursion depth is $\mathcal{O}(\log\log n)$. Thus, if the expected computational complexity of the partition is $\mathcal{O}(n)$ and the probability that the bucketing “fails” is sufficiently small, the total computational complexity is $\mathcal{O}(n\log\log n)$. Under mild assumptions on the distribution, we can prove that this fast and accurate partition can be achieved by using the Piecewise Constant Function (PCF) as the CDF model.
@@ 3.1 Method Overview
Let $\mathcal{D}~{}(\subset{\mathbb{R}})$ be the range of possible values of the input array and $\bm{x}~{}(\in\mathcal{D}^{n})$ be the input array. If the length of the input array, denoted as $|\bm{x}|$, is less than $\tau$, our algorithm sorts the input array using a “standard” sorting method with $\mathcal{O}(n\log n)$ complexity, such as IntroSort, where $\tau$ is a predetermined constant. Otherwise, if $|\bm{x}|\geq\tau$, model-based bucketing is performed.
@@
The model-based bucketing method $\mathcal{M}$ takes an input array $\bm{x}$ and partitions it into several buckets. First, it determines the parameter $\gamma~{}(\in\mathbb{N})$, which determines the number of buckets. $\gamma$ is determined as a function of the array length $n$. The number of buckets returned by $\mathcal{M}$ is $\gamma+1$. Next, all or some elements of $\bm{x}$ are used to train the CDF model $\tilde{F}\colon\mathcal{D}\mapsto[0,1]$. The $\tilde{F}(q)$ is trained to predict the empirical cumulative distribution function of $\bm{x}$ defined as:
$F_{\bm{x}}(q)\coloneqq\frac{|\{i\in\{1,\dots,n\}\mid x_{i}\leq q\}|}{|\bm{x}|}.$ (4)
@@
The specific model and training method is explained in Section 3.2. Finally, the CDF model is used to partition the input array $\bm{x}$ into $\gamma+1$ buckets. All $\gamma+1$ buckets, $\{\bm{c}_{j}\}_{j=1}^{\gamma+1}$, are initialized to be empty, and then for each $i\in\{1,\dots,n\}$, $x_{i}$ is appended to $\bm{c}_{\left\lfloor\tilde{F}(x_{i})\gamma\right\rfloor+1}$. This is based on the intuition that the number of elements less than or equal to $x_{i}$ in the array $\bm{x}$ (i.e., |$|\{j\in\{1,\dots,n\}\mid x_{j}\leq x_{i}\}|$) is approximately equal to $n\tilde{F}(x_{i})$.
The CDF model $\tilde{F}$ must be a non-decreasing function to guarantee the computational complexity of our algorithm. In this case, the bucket with the larger ID gets the larger value, i.e.,
$p\in\bm{c}_{j},q\in\bm{c}_{k},1\leq j<k\leq\gamma+1\Rightarrow p<q.$ (5)
@@
This means that each bucket is responsible for a disjoint and continuous interval. Let $t_{j}=\min_{x\in\bm{c}_{j}}x~{}(j=1,\dots,\gamma+1),~{}t_{\gamma+2}=\infty$, then the $j$-th bucket $\bm{c}_{j}$ ($j=1,\dots,\gamma+1$) is responsible for a continuous interval $\mathcal{I}_{j}\coloneqq[t_{j},t_{j+1})$.
@@
After model-based bucketing, our algorithm determines for each bucket whether the bucketing was “successful” or not. For each $j=1,\dots,\gamma+1$, we check whether the size of bucket $\bm{c}_{j}$ is less than $\delta$, where $\delta$ is an integer determined by $n$. If $|\bm{c}_{j}|\geq\delta$ (which means the bucketing “fails”), the bucket is sorted using the “standard” sorting method (e.g., IntroSort). If $|\bm{c}_{j}|<\delta$ (which means the bucketing “succeeds”), the bucket is sorted by recursively calling our Learned Sort algorithm. Note that the parameters such as $\gamma$ and $\delta$ are redetermined according to the size of the bucket (i.e., the input array in the next recursion step), and the CDF model is retrained for each bucket. After each bucket is sorted, the buckets are concatenated. The input array $\bm{x}$ is sorted by the above procedure. The basic idea of our algorithm is visualized in Figure 1, and the pseudocode is given in Algorithm 1.
@@
The following is a lemma about the worst-case complexity of our Learnd Sort as defined above.
@@
Assume that there exists an model-based bucketing algorithm $\mathcal{M}$ such that $\mathcal{M}$ can perform bucketing (including model training and inferences) an array of length $n$ into $\gamma+1=\mathcal{O}(n)$ buckets with a worst-case complexity of $\mathcal{O}(n)$. Also, assume that the “standard” sort algorithm used for the too large or too small buckets has a worst-case complexity of $\mathcal{O}(n\log n)$. The worst-case complexity of our Learned Sort with such $\mathcal{M}$ and $\delta=\lfloor n^{d}\rfloor$ (where $d$ is a constant satisfying $10<d<$ 1) is $\mathcal{O}(n\log n)$.
This lemma can be intuitively shown from the following two points: (i) the maximum recursion depth is $\mathcal{O}(\log\log n)$, and (ii) each element of the input array $\bm{x}$ undergoes several bucketing and only one “standard” sort. (i) can be shown from the fact that the size of the bucket in the $i$-th recursion depth is less than $n^{d^{i}}$, and (ii) is evident from the algorithm’s design since the buckets sorted by “standard” sort are now left only to be concatenated. The exact proof is given in Section A.1.
Next, we introduce an important lemma about the expected complexity of our Learned Sort. The following assumption is necessary to guarantee the expected complexity.
@@
The input array $\bm{x}\in\mathcal{D}^{n}$ is formed by independent sampling according to a probability density distribution $f(x)\colon\mathcal{D}\mapsto\mathbb{R}_{\geq 0}$.
We define $f_{\mathcal{I}}(x)\colon\mathcal{I}\mapsto\mathbb{R}_{\geq 0}$ to be the conditional probability density distribution of $f(x)$ under the condition that $x\in\mathcal{I}$ for a interval $\mathcal{I}\subset\mathcal{D}$, i.e.,
[…]

### v3

@@ Abstract
Sorting is one of the most fundamental algorithms in computer science. Recently, Learned Sorts, which use machine learning to improve sorting speed, have attracted attention. While existing studies show that Learned Sort is empirically faster than classical sorting algorithms, they do not provide theoretical guarantees about its computational complexity. We propose Piecewise Constant Function (PCF) Learned Sort, a theoretically guaranteed Learned Sort algorithm. We prove that the expected complexity of PCF Learned Sort is $\mathcal{O}(n\log\log n)$ under mild assumptions on the data distribution. We also confirm empirically that PCF Learned Sort has a computational complexity of $\mathcal{O}(n\log\log n)$ on both synthetic and real datasets. This is the first study to theoretically support the empirical success of Learned Sort, and provides evidence for why Learned Sort is fast.
@@ 1 Introduction
Sorting is one of the most fundamental algorithms in computer science and has been extensively studied for many years. Recently, a novel sorting method called Learned Sort has been proposed (kraska2021sagedb). In Learned Sort, a machine learning model is trained to estimate the distribution of elements in the input array, specifically, the cumulative distribution function (CDF). The model’s predictions are then used to rearrange the elements, followed by a minor refinement step to complete the sorting process. Empirical results show that Learned Sort is faster than classical sorting algorithms, including highly optimized counting-based sorting algorithms, comparison sorting algorithms, and hybrid sorting algorithms.
On the other hand, there are few theoretical guarantees regarding the computational complexity of Learned Sort. The first proposed Learned Sort algorithm (kraska2021sagedb) has a best-case complexity of $\mathcal{O}(n)$, but its expected or worst-case complexity is not discussed. The more efficient Learned Sort algorithms proposed later (kristo2020case; kristo2021defeating) also have $\mathcal{O}(n)$ best-case complexity, but $\mathcal{O}(n^{2})$ worst-case complexity (or $\mathcal{O}(n\log n)$ with some modifications). The goal of this paper is to develop a Learned Sort that is theoretically guaranteed to be computationally efficient.
We propose Piecewise Constant Function (PCF) Learned Sort, which can sort with an expected complexity $\mathcal{O}(n\log\log n)$ under mild assumptions on the data distribution. In addition, we show that PCF Learned Sort admits a worst-case complexity guarantee that depends on the choice of the internal sorting algorithm. For instance, if MergeSort—whose worst-case complexity is $\mathcal{O}(n\log n)$—is used, then the worst-case complexity of PCF Learned Sort is $\mathcal{O}(n\log n)$. If we instead use the algorithm by han2020sorting, which has a worst-case complexity of $\mathcal{O}(n\sqrt{\log n})$, then the worst-case complexity of PCF Learned Sort becomes $\mathcal{O}(n\sqrt{\log n})$. We then empirically confirm that our Learned Sort can sort with a complexity of $\mathcal{O}(n\log\log n)$ on both synthetic and real datasets.
Independently and concurrently, zeighami2024theoretical explored complexity-guaranteed Learned Sort. While our PCF Learned Sort is motivated by similar principles and incorporates comparable design choices, several key distinctions exist between our approach and theirs. A comprehensive comparison between our method and that of zeighami2024theoretical is provided in Section˜5.
This paper is organized as follows. Section˜2 reviews related literature. Section˜3 introduces PCF Learned Sort and provides complexity theorems with proof sketches. Section˜4 provides empirical validation of the theoretical results. We discuss implications and limitations in Section˜5, and conclude in Section˜6.
@@ 2 Related Work
Our research is in the context of algorithms with machine learning (Section˜2.1). There are two types of sorting algorithms, comparison sorts (Section˜2.2) and non-comparison sorts (Section˜2.3), and our proposed method is a non-comparison sort. However, the idea, implementation, and proof of computational complexity of our method are similar to those of sample sort, which is a type of comparison sort. Furthermore, our proposed algorithm and proof of computational complexity are based on those of Learned Index (Section˜2.4).
2.1 Algorithms with Machine Learning
Our work lies in the emerging area of Learned Sort, where machine learning models accelerate sorting. The first approach by kraska2021sagedb trained a model $\tilde{F}(q)$ to approximate the CDF $F(q)$ and placed each element at position $n\tilde{F}(x)$, and then applies insertion sort to complete the ordering. While this yields $\mathcal{O}(n)$ in favorable cases, the insertion-based refinement causes a $\mathcal{O}(n^{2})$ worst case. Later implementations improved cache efficiency (kristo2020case), robustness to duplicates (kristo2021defeating), and even introduced cache-friendly CDF models tailored for sorting (ferragina2025balancedsort), but still relied on insertion sort and thus retained a quadratic worst-case complexity. Using stronger fallback algorithms such as Introsort (musser1997introspective) or TimSort (mcilroy1993optimistic) reduces this to $\mathcal{O}(n\log n)$, though this only matches classical comparison-based sorts and does not fully explain the empirical advantage of Learned Sort.
Parallelization has also been explored: carvalho2022towards proposed IPLS, integrating learned partitioning into IPS4o (axtmann2017), and carvalho2023learnedsort framed LearnedSort as a sample sort with a parallel IPS4o implementation. These studies primarily highlight the engineering potential of learned sorting in parallel settings, which is complementary but orthogonal to our focus on providing theoretical guarantees for the sequential case.
In a broader context, li2020strong showed that neural programs can achieve strong generalization and replicate efficient algorithmic behaviors such as sorting, demonstrating that machine learning can recover algorithmic complexity classes like $\mathcal{O}(n\log n)$ without explicit manual design. This provides a useful perspective on the generalizability of learned algorithms.
Related to this, the field of algorithms with predictions studies how machine learning predictions accelerate classical algorithms (mitzenmacher2022algorithms), with applications in caching (narayanan2018deepcache; rohatgi2020near; lykouris2021competitive; im2022parsimonious), ski rental (purohit2018improving; gollapudi2019online; shin2023improved), scheduling (gollapudi2019online; lattanzi2020online; lassota2023minimalistic), and matching (antoniadis2023online; dinitz2021faster; sakaue2022discrete). Sorting with predictions has also been analyzed (lu2021generalized; chan2023generalized; erlebach2023sorting), including tight guarantees by bai2023sorting. In the context of algorithms with predictions, machine learning predictions are typically assumed to be available at no cost, and the models are treated as an opaque box. This exclusion contrasts with our problem setting, where we ensure that the computational complexity covers the entire process from receiving an unsorted array to returning a sorted array.
@@ 2.2 Comparison Sorts
Sorting algorithms that use comparisons between keys and require no other information about the keys are called comparison sorts. It is well-known that the worst-case complexity of a comparison sort is at least $\Omega(n\log n)$. Commonly used comparison sorting algorithms include QuickSort, heap sort, merge sort, and insertion sort. The GNU Standard Template Library in C++ uses Introsort (musser1997introspective), an algorithm that combines QuickSort, heap sort, and insertion sort. Java (javalistsort2023) and Python up to version 3.10 (python_timsort2002) use TimSort (mcilroy1993optimistic), an improved version of merge sort. Python 3.11 and later use Powersort (munro2018nearly), a merge sort that determines a near-optimal merge order.
Sample sort (frazer1970samplesort) extends QuickSort by using multiple pivots instead of one. Sample sort samples a small number of keys from the array, determines multiple pivots, and uses them to partition the array into multiple buckets. The partitioning is repeated recursively until the array is sufficiently small. Among its implementations, the in-place parallel superscalar sample sort (axtmann2017) was introduced as a highly efficient parallel algorithm, and later engineering refinements further improved its performance (axtmann2022engineering). Its computational and cache efficiency is theoretically guaranteed by a theorem about the probability that the pivots partition the array (nearly) equally.
2.3 Non-Comparison Sorts
Non-comparison sorts use information other than key comparison. Radix sort and counting sort are the most common types of non-comparison sorts. Radix sort uses counting sort as a subroutine for each digit. When the number of digits in the array element is $w$, the computational complexity of radix sort is $\mathcal{O}(wn)$. Thus, radix sort is particularly effective when the number of digits is small. There are several variants of radix sort, such as Spreadsort (ross2002spreadsort), which integrates the advantages of comparison sort into radix sort and is implemented in the Boost C++ Libraries, and RegionSort (obeya2019theoretically), which enables efficient parallelization by modeling and resolving dependencies among the element swaps.
In addition, non-comparison sorting algorithms tailored for specific data types have been developed. For integer arrays, a deterministic algorithm with worst-case complexity of $\mathcal{O}(n\log\log n)$ (han2002deterministic) and a randomized algorithm with expected complexity of $\mathcal{O}(n\sqrt{\log\log n})$ (han2002integer) have been proposed. For real-valued arrays, recent advances have led to the development of a sorting algorithm with a worst-case complexity of $\mathcal{O}(n\sqrt{\log n})$ (han2020sorting). Our PCF Learned Sort also targets real-valued arrays and, under mild assumptions on the distribution, achieves an expected complexity of $\mathcal{O}(n\log\log n)$. Moreover, when the algorithm of han2020sorting is used as the internal sorting component, PCF Learned Sort has a worst-case complexity of $\mathcal{O}(n\sqrt{\log n})$, which matches that of han2020sorting.
@@ 2.4 Learned Index
kraska2018case_li showed that index data structures such as B-trees and Bloom filters can be made faster or more memory efficient by combining them with machine learning models and named such novel data structures Learned Index. Since then, learning augmented B-trees (wang2020sindex; kipf2020radixspline; ferragina2020pgm; zeighami2023distribution), Bloom filters (mitzenmacher2018model; vaidya2020partitioned; sato2023fast; sato2024fast), and even range-minimum query structures (ferragina2025fl) have been proposed. There are several works on learning augmented B-trees whose performance is theoretically guaranteed. PGM-index (ferragina2020pgm) is a learning augmented B-tree that is guaranteed to have the same worst-case query complexity as the classical B-tree, i.e., $\mathcal{O}(\log n)$. zeighami2023distribution proposed a learning augmented B-tree with an expected query complexity of $\mathcal{O}(\log\log n)$ under mild assumptions on the distribution. More broadly, this line of research connects to earlier studies on compressed data structures with provable guarantees, such as the foundational work of sadakane2007compressed.
@@ 3 Methods
This section introduces our Learned Sort framework and its PCF-based instantiation. In Section˜3.1, we set up the notation, in Section˜3.2 we present the general framework, and in Section˜3.3 we instantiate it with a PCF-based CDF model to obtain concrete complexity guarantees.
3.1 Notation and Setup
Following the learned-sort paradigm kraska2021sagedb; kristo2020case; kristo2021defeating, our learned sort algorithm recursively applies model-based bucketing. Model-based bucketing partitions the array into multiple buckets using a CDF model trained on the input so that the bucket with the larger ID gets the larger value.
Let $\mathcal{D}\subseteq\mathbb{R}$ be the domain of possible key values and $\bm{x}\in\mathcal{D}^{n}$ be the input array. For vector/array $\bm{v}$, we write $|\bm{v}|$ for its length (number of elements), e.g., $|\bm{x}|=n$. If $|\bm{x}|$ is smaller than a fixed threshold $\tau\in\mathbb{N}$, the algorithm falls back to a classic sorting algorithm. We refer to this fallback classical sorting algorithm as the standard sort. As the standard sort, we can use any sorting algorithm, such as IntroSort, QuickSort, or the algorithm of han2020sorting.
For model-based bucketing, we define functions $\alpha(n),\beta(n),\gamma(n),\delta(n):\mathbb{N}\to\mathbb{N}$ as follows: $\alpha(n)$ is the number of samples used to train the CDF model, $\beta(n)$ is the number of PCF bins (a hyperparameter; see Section 3.3), $\gamma(n)+1$ is the number of buckets, and $\delta(n)$ is the threshold that determines the behavior after bucketing (see Section 3.2). For brevity, we write $\alpha,\beta,\gamma,\delta$.
3.2 Learned-Sort Framework
Figure 1: PCF Learned Sort: First, the input array is partitioned into $\gamma+1$ buckets using a CDF model-based method. Buckets larger than $\delta$ or smaller than $\tau$ are sorted with a standard sort (e.g., IntroSort or QuickSort). Otherwise, the recursive model-based bucketing is repeated. Finally, the sorted arrays are concatenated. The CDF model used for bucketing is a Piecewise Constant Function (PCF). The function is constant within each interval, and the interval widths are constant.
@@ 3.2 Learned-Sort Framework
1:Input: $\bm{x}\in\mathbb{R}^{n}$Output: $\bm{x}_{\mathrm{sorted}}\in\mathbb{R}^{n}$ (the sorted version of $\bm{x}$)
2:Subroutines:
@@
5:function Learned-Sort($\bm{x}$)
6:  $n\leftarrow|\bm{x}|$
7:  if $n<\tau$ then $\triangleright$ Small bucket
@@
9:  $\tilde{F}(q)\leftarrow\textsc{CDF-Model}(\bm{x})$ $\triangleright$ Model-based bucketing
10:  $\bm{c}_{1}\leftarrow[~],~\bm{c}_{2}\leftarrow[~],~\dots~,~\bm{c}_{\gamma+1}\leftarrow[~]$
11:  for $i=1,2,\dots,n$
12:   $j\leftarrow\left\lfloor\tilde{F}(x_{i})\gamma\right\rfloor+1$
13:   $\bm{c}_{j}.\textsc{Append}(x_{i})$
14:  for $j=1,2,\dots,\gamma+1$ $\triangleright$ Recursively sort and concatenate
15:   if $|\bm{c}_{j}|\geq\delta$ then
16:     $\bm{c}_{j}\leftarrow\textsc{Standard-Sort}(\bm{c}_{j})$
17:   else
18:     $\bm{c}_{j}\leftarrow\textsc{Learned-Sort}(\bm{c}_{j})$
19:  $\bm{x}_{\mathrm{sorted}}\leftarrow\textsc{Concatenate}(\bm{c}_{1},\bm{c}_{2},\dots,\bm{c}_{\gamma+1})$
20:  return $\bm{x}_{\mathrm{sorted}}$
Algorithm description.
At a high level, our algorithm is a recursive model-based bucketing scheme augmented with an exception-handling mechanism for bucketing “failures.” This exception-handling mechanism greatly simplifies the analysis of both expected and worst-case complexity. The overall workflow is visualized in Figure˜1, and the pseudocode is given in Algorithm˜1.
@@
If the length of the input array is less than $\tau$, our algorithm sorts the input array using a standard sort (e.g., IntroSort or QuickSort). Otherwise, model-based bucketing is performed.
@@
The model-based bucketing method $\mathcal{M}$ takes an input array $\bm{x}$ and partitions it into several buckets. First, all or some elements of $\bm{x}$ are used to train the CDF model $\tilde{F}\colon\mathcal{D}\to[0,1]$. The $\tilde{F}(q)$ is trained to approximate the empirical CDF $F_{\bm{x}}(q)=|{i\in\{1,\dots,n\}\mid x_{i}\leq q}|/|\bm{x}|$. Any non-decreasing model can serve as the CDF model $\tilde{F}(q)$ (e.g., linear models, monotonic MLPs, or the PCF introduced in Section˜3.3).
@@
After training, the CDF model is used to partition the input array $\bm{x}$ into $\gamma+1$ buckets. All $\gamma+1$ buckets, $\{\bm{c}_{j}\}_{j=1}^{\gamma+1}$, are initialized to be empty, and then for each $i\in\{1,\dots,n\}$, $x_{i}$ is appended to $\bm{c}_{\left\lfloor\tilde{F}(x_{i})\gamma\right\rfloor+1}$. This is based on the intuition that the number of elements less than or equal to $x_{i}$ in the array $\bm{x}$ (i.e., $|\{j\in\{1,\dots,n\}\mid x_{j}\leq x_{i}\}|$) is approximately equal to $n\tilde{F}(x_{i})$.
@@
We restrict the CDF model $\tilde{F}$ to non-decreasing functions to ensure that the bucket with the larger ID gets the larger value, i.e., $p\in\bm{c}_{j}\;\land\;q\in\bm{c}_{k}\;\land\;j<k\;\;\Rightarrow\;\;p<q$. This means that each bucket is responsible for a disjoint and continuous interval. Let $t_{j}=\min_{x\in\bm{c}_{j}}x~(j=1,\dots,\gamma+1),~t_{\gamma+2}=\infty$, then the $j$-th bucket $\bm{c}_{j}$ ($j=1,\dots,\gamma+1$) is responsible for a continuous interval $\mathcal{I}_{j}\coloneqq[t_{j},t_{j+1})$.
@@
After model-based bucketing, our algorithm determines for each bucket whether the bucketing “succeeds” or “fails.” For each $j\in\{1,\dots,\gamma+1\}$, we check whether the size of bucket $\bm{c}_{j}$ is less than $\delta$. If $|\bm{c}_{j}|\geq\delta$ (which means the bucketing “fails”), the bucket is sorted using the standard sort (e.g., IntroSort or QuickSort). If $|\bm{c}_{j}|<\delta$ (which means the bucketing “succeeds”), the bucket is sorted by recursively calling our Learned Sort algorithm. Note that the parameters such as $\gamma$ and $\delta$ are redetermined for each recursion according to the size of the bucket (i.e., the input array in the next recursion step), and the CDF model is retrained for each bucket. Finally, the sorted buckets are concatenated to form the sorted array.
Worst-case complexity.
@@
The following lemma guarantees the worst-case complexity of our Learned Sort.
@@
Assume that there exists a model-based bucketing algorithm $\mathcal{M}$ such that $\mathcal{M}$ can perform bucketing (including model training and inferences) an array of length $n$ into $\gamma+1=\mathcal{O}(n)$ buckets with a worst-case complexity of $\mathcal{O}(n)$. Also, assume that the standard sort has a worst-case complexity of $\mathcal{O}(nU(n))$, where $U(n)$ is a non-decreasing function. Then, the worst-case complexity of our Learned Sort with such $\mathcal{M}$ and $\delta=\lfloor n^{d}\rfloor$ (where $d$ is a constant satisfying $0<d<1$) is $\mathcal{O}(nU(n)+n\log\log n)$.
This lemma can be intuitively shown from the following two points: (1) the maximum recursion depth is $\mathcal{O}(\log\log n)$, and (2) each element of the input array $\bm{x}$ undergoes several bucketing and only one standard sort. The first point can be shown from the fact that the size of the bucket in the $i$-th recursion depth is less than $n^{d^{i}}$. The second point is evident from the algorithm’s design since the buckets sorted by standard sort are now left only to be concatenated. The exact proof is given in Section˜A.1.
Note that this guarantee critically relies on the exception-handling mechanism: without it, the recursion depth could reach $\Theta(n)$ and the worst-case time would degrade to $\Theta(n^{2})$, whereas with it the recursion depth is always bounded by $O(\log\log n)$.
Expected complexity.
Next, we introduce a lemma about the expected computational complexity of our Learned Sort. The following assumption is necessary to guarantee the expected computational complexity.
@@
The input array $\bm{x}\in\mathcal{D}^{n}$ is formed by independent sampling according to a probability density distribution $f(x)\colon\mathcal{D}\to\mathbb{R}_{\geq 0}$.
We define $f_{\mathcal{I}}(x)\colon\mathcal{I}\to\mathbb{R}_{\geq 0}$ to be the conditional probability density distribution of $f(x)$ under the condition that $x\in\mathcal{I}$ for a interval $\mathcal{I}\subseteq\mathcal{D}$, i.e., $f_{\mathcal{I}}(x)\coloneqq\frac{f(x)}{\int_{\mathcal{I}}f(y)dy}$.